# MagicPose4D: Crafting Articulated Models with Appearance and Motion Control

### v1

@@ 4.1 Qualitative and Quantitative Comparison
Our Framework v.s. Existing 4D Generation. Text/image-prompts-based 4D generation has been a popular trend. To compare with a representative recent work Animate124 Zhao et al. (2023), we feed the reference image, which provides the identity, and text prompt, which describes the motion of the generated object, into their model and optimize the NeRF with SDS-loss. In our case, we use the reference image to generate the target mesh with an Image-to-3D model and transfer the motion of the reference mesh sequence. We visualize the rendered video from different viewpoints from Animate124 and compare it with MagicPose4D in Fig. 6(a) and Fig. 6(b). More video comparisons are in the supplementary material. MagicPose4D provides more temporally consistent and smooth generation with the learning of motion deformation. Since there are no ground truth mesh sequences to evaluate the generation, we provide a comprehensive user study for Motion Transfering and 4D Generation for a qualitative evaluation in Sec. 6.4 and Tab.3 to conclude our findings.
@@
Table 1: 2D Keypoint Transfer Accuracy. This table presents the 2D keypoint transfer accuracy on DAVIS and PlanetZoo datasets. For classes from DAVIS (camel) and PlanetZoo (giraffe, tiger), we carry out multiple executions and report the mean of the observed accuracy in each case. Note that the results for LASR and ViSER here are from our executions of the codes provided by the authors.
@@
ViSER 71.7 73.7 65.1 72.7 78.3 51.2 68.4 68.9 60.3 55.7 57.3 65.8
LASR 75.2 80.3 60.3 83.1 55.3 56.3 70.4 69.5 63.1 57.4 60.3 66.5
MagicPose4D 78.9 83.1 67.9 86.2 84.3 62.9 73.3 70.6 67.1 60.2 61.9 72.4
Figure 7: Ablation Experiments.
Table 2: User study of MagicPose4D on Motion Transfer. We collect the rating results from 50 participants of eleven mesh-sequence comparison experiments. The scale of rating is 0 (low) - 5 (high). The participants found that MagicPose4D generates the best motion transfer results. Criteria for judgment: 1) The generated motion should match the reference pose mesh sequence. 2) The identity of the transferred mesh should match the identity reference. 3) The generated mesh sequence should be consistent and smooth. More details can be found in Appendix Sec.6.4
@@
MagicPose4D 4.92 4.47 4.69 4.39 4.45 4.45 3.90 4.12 4.02 4.10 4.20 4.34
@@
Table 3: User study of MagicPose4D on 4D Generation compared to Animate124 Zhao et al. (2023). The settings are the same as the study on Motion Transfer. Criteria for judgment: 1) The generation’s identity should match the reference image. 2) The video should be temporal-consistent and smooth.
@@
4.2 Diagnostic
Effectiveness of Global-Local Chamfer Loss. Relying solely on the geometric information contained in a short monocular video often makes it challenging to achieve satisfactory 3D reconstruction results. This difficulty arises because most videos typically offer sparse viewpoints and the presence of non-rigid object deformations increases the optimization space. Consequently, many methods that rely only on 2D supervision Yang et al. (2021b, 2022b); Wu et al. (2023) struggle to ensure effectiveness on in-the-wild videos. To enhance the generalization capability of the reconstruction module, we utilize existing image-to-3D model methods to obtain geometric priors. By using these models, we generate pseudo-3D meshes for each frame as 3D supervision. We employ the Chamfer distance as the 3D loss objective to align the predicted 3D mesh closely with the pseudo-3D mesh in terms of point distribution. However, global Chamfer distance alone can only ensure similar point distributions between the two meshes and does not guarantee the correspondence of points as expected. For instance, as shown in Fig.7(a), without using the local Chamfer loss, the overall point distribution might be similar, but the mesh shape could be entirely incorrect, such as the left and right legs being swapped or vertices originally on the right front leg being moved near the left hind leg. By incorporating the local Chamfer loss, we enforce consistency between each part of the predicted mesh and the pseudo mesh, thus ensuring pose consistency and significantly mitigating the previously mentioned issues.
5 Conclusion & Limitations
The main limitations of our method and existing works are: (i) Deformation based on blend skinning relies on accurate and robust skeletons and skinning weights predictions, facing a trade-off between generalization and accuracy. Learning-based methods have limited generalization due to restricted training datasets, while non-learning methods suffer from inductive bias, leading to suboptimal results. (ii) MagicPose4D can infer poses quickly for pose transfer without training, but 4D reconstruction requires significant training (10 hours on an L40S). (iii) Our method struggles with detailed motion control, such as fingers and facial features, due to the challenge of capturing fine-grain details during 4D reconstruction. These issues represent future research directions.

### v2

@@ 4 Experiments
To compare with representative recent works 4Dfy Bahmani et al. (2024b) and Animate124 Zhao et al. (2023), we feed the reference image, which provides the identity, and text prompt, which describes the motion of the generated object (4Dfy only uses text prompts), into their model and optimize the NeRF with SDS-loss. To compare with Stag4D, we first leverage SVD Blattmann et al. (2023b) and DMT Yatim et al. (2024) to generate video according to the input reference text or video and then use Stag4D Zeng et al. (2024) to do reconstruction from the generated video. In our case, we use the reference image to generate the target mesh with an Image-to-3D model and transfer the motion of the reference mesh sequence, which is built by our 4D reconstruction module. We visualize the rendered video from different viewpoints from those SOTA methods and compare them with MagicPose4D in Fig. 5(a) and Fig. 8. More video comparisons are in the supplementary material and on our web page. MagicPose4D provides more temporally consistent and smooth generation with the learning of motion deformation. Since there are no ground truth mesh sequences to evaluate the generation, we provide a comprehensive user study for 4D Generation for a qualitative evaluation in Sec. A.5 and Tab.1 and 6 to conclude our findings.
4.2 4D Cross-Category Motion Transfer
MagicPose4D is able to learn and retarget the motion from dynamic mesh prompts. In this section, we present motion transfer results from MagicPose4D in Fig. 4. The target identities and reference motion sequences can either be humanoid or animal subjects.
We further compare the motion transfer ability for mesh generation quality with recent methods 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023) in Fig. 5 (c) and Tab.3. Both methods use a deep neural network to learn latent shape codes to retarget the motions and are trained/evaluated on SMPL Loper et al. (2023b)(humanoid) and SMAL Zuffi et al. (2017)(animal). Since they do not include disentangled components or shape deformation modules to represent the structural information of 3D meshes, these methods rely heavily on pre-processed training data with high-quality mesh annotations. Due to these constraints, these baselines cannot generalize well to in-the-wild target identities or reference motions. In contrast, MagicPose4D generates temporally consistent and smooth motions, while strictly preserving the identity and appearance of the target mesh. It is also worth noting that because previous methods focus only on pose transfer without considering the motion trajectory, the generated object always stays in the same position. This can be observed in side-by-side video comparisons on our anonymous webpage: https://magicpose4d.github.io/.
4.3 4D Reconstruction
Since the primary goal of our 4D reconstruction module is to better capture motion, we focus on the accuracy of keypoint reconstruction. Therefore, we use keypoint transfer accuracy as the quantitative evaluation metric. We examine the performance of MagicPose4D against the SOTA methods, S3O Zhang et al. (2024b), LIMR Zhang et al. (2024c), LASR Yang et al. (2021a) and ViSER Yang et al. (2021c). As shown in Tab.2, MagicPose4D consistently outperforms them for all animal subjects with a notable margin. When we run the author-provided codes for LASR, and ViSER on DAVIS, and PlanetZoo, we find that their results are highly variable across different runs. The results we report for the mean accuracies we have obtained over multiple runs ($5$) of all methods. Furthermore, as depicted in Fig.7, we compare the results of MagicPose4D with LASR Yang et al. (2021a) and BANMo Yang et al. (2022a). These methods fall short of achieving ideal reconstruction, due to a lack of structural information about the objects, since they learn from only a single monocular video, which contains sparse views. NeRF-based methods such as BANMo Yang et al. (2022b) and MagicPony Wu et al. (2023) require particularly dense views from multiple cameras or long videos to achieve decent results, while ours achieves good reconstructions from only sparse views.
Figure 5: 4D Generation. a) and b): Comparison of MagicPose4D with Animate124 Zhao et al. (2023), Motion Transfer. c): Comparison of MagicPose4D with 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023). Videos are in Sec.A.2.
@@
Table 2: 2D Keypoint Transfer Accuracy. This table presents the 2D keypoint transfer accuracy on DAVIS and PlanetZoo datasets. For all methods, we carry out multiple executions and report the mean of the observed accuracy in each case. Note that all experiment results of MP4D are without the skeleton templates and learning underlying structure following Zhang et al. (2024b).
@@
ViSER Yang et al. (2021c) 71.7 73.7 65.1 72.7 78.3 51.2 68.4 68.9 60.3 55.7 57.3 65.8
LASR Yang et al. (2021b) 75.2 80.3 60.3 83.1 55.3 56.3 70.4 69.5 63.1 57.4 60.3 66.5
LIMR Zhang et al. (2024c) 77.5 79.6 65.4 84.7 79.6 59.3 64.9 68.3 64.1 57.1 55.2 68.7
S3O Zhang et al. (2024b) 79.8 83.2 67.3 86.7 84.7 63.1 72.8 71.3 66.4 61.5 62.1 72.6
MP4D (Ours) 80.4 84.8 69.7 87.9 86.0 64.7 75.2 72.4 68.9 61.9 63.7 74.2
Figure 6: Analysis. a): Visualization for the effectiveness of Local Chamfer Distance Loss. b): Performing pose transfer independently for each frame results in poor temporal consistency. c) View 2 is a failure case generated by imaged-to-3D models due to self-occlusion.
Table 3: User study of MagicPose4D on Motion Transfer. We collect the rating results from 50 participants of eleven mesh-sequence comparison experiments. The scale of rating is 0 (low) - 5 (high). Criteria for judgment: 1) The generated motion should match the reference pose mesh sequence. 2) The identity of the transferred mesh should match the identity reference. 3) The generated mesh sequence should be consistent and smooth. More details can be found in Appendix Sec.A.5
@@
MP4D(Ours) 4.92 4.47 4.69 4.39 4.45 4.45 3.90 4.12 4.02 4.10 4.20 4.34
4.4 Diagnostic
Effectiveness of Global-Local Chamfer Loss. Relying solely on the geometric information contained in a short monocular video often makes it challenging to achieve satisfactory 3D reconstruction results. This difficulty arises because most videos typically offer sparse viewpoints and the presence of non-rigid object deformations increases the optimization space. Consequently, many methods that rely only on 2D supervision Yang et al. (2021b; 2022b); Wu et al. (2023) struggle to ensure effectiveness on in-the-wild videos. To enhance the generalization capability of the reconstruction module, we utilize existing image-to-3D model methods to obtain geometric priors. By using these models, we generate pseudo-3D meshes for each frame as 3D supervision. We employ the Chamfer distance as the 3D loss objective to align the predicted 3D mesh closely with the pseudo-3D mesh in terms of point distribution. However, global Chamfer distance alone can only ensure similar point distributions between the two meshes and does not guarantee the correspondence of points as expected. For instance, as shown in Fig.6(a), without using the local Chamfer loss, the overall point distribution might be similar, but the mesh shape could be entirely incorrect, such as the left and right legs being swapped or vertices originally on the right front leg being moved near the left hind leg. By incorporating the local Chamfer loss, we enforce consistency between each part of the predicted mesh and the pseudo mesh, thus ensuring pose consistency and significantly mitigating the previously mentioned issues.
5 Conclusion & Limitations
The main limitations of our method and existing works are: (i) Deformation based on blend skinning relies on accurate and robust skeletons and skinning weights predictions, facing a trade-off between generalization and accuracy. Learning-based methods have limited generalization due to restricted training datasets, while non-learning methods suffer from inductive bias, leading to suboptimal results. (ii) MagicPose4D can infer poses quickly for pose transfer without training, but 4D reconstruction requires significant training (5 hours on an L40S). (iii) Our method struggles with detailed motion control, such as fingers and facial features, due to the challenge of capturing fine-grain details during 4D reconstruction. These issues represent future research directions.
6 Acknowledgments
This work was supported by the Office of Naval Research (grant N00014-20-1-2444) and the USDA National Institute of Food and Agriculture (grant 2020-67021-32799 /1024178). Support for computational resources was provided by the National Science Foundation (awardOAC 2005572) and the State of Illinois, through Delta Gropp et al. (2023), a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
Appendix A Appendix
In this section, we aim to provide additional information not included in the main text due to length constraints. This supplemental content includes:
  1. 1.
    Support prompts comparison with existing 4D generation methods in Tab.4.
  2. 2.
    Key Differences Between the Two Stages in 4D Reconstruction Module in Sec.A.1.
  3. 3.
    Video results (Sec.A.2)
  4. 4.
    Implementation details (Sec.A.4)
  5. 5.
    Information about the dataset used for evaluation (Sec.A.3)
  6. 6.
    In-depth explanation of loss and regularization terms (Sec.A.6)
  7. 7.
    Further experimental results: (i) comparison with existing 4D reconstruction methods in Fig.7, 4D generation method in Fig.8 and Tab.6, (ii) comparison with existing pose transfer methods in Fig.9, (iii) skeleton extraction without skeleton template in Fig.10, and (iv) skeleton and skinning weights results in Fig.11.
  8. 8.
    Details of user study (Sec.A.5)
  9. 9.
    Broader social impacts (Sec.B)
Table 4: Support Prompts Comparison. Random represents no control of motion.
Method Appearance Motion
Text Image Text Video Trajectory Random
Animate124 Zhao et al. (2023)
4D-fy  Bahmani et al. (2023)
DreamGaussian4D Ren et al. (2023)
AYG Ling et al. (2023)
Dream-in-4D Zheng et al. (2023)
TC4D Bahmani et al. (2024a)
MagicPose4D
Figure 8: 4D Generation. Comparison of MagicPose4D with 4Dfy, Stag4D, DG4D, DMT+Stag4D, and Animate124.
Figure 9: Motion Transfer. Comparison with 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023)
Figure 10: Generated skeleton result without skeleton template. Here we visualize two different poses for wolf and bear from two different views.
Figure 11: Skinning Weights and Skeleton Results from MagicPose4D. The skeletons can fit to both humanoid and animal objects.
Figure 12: Post Texture Editing.
A.1 Key Differences Between the Two Stages in 4D Reconstruction
In the first stage of reconstruction, similar to many existing methods [38, 31, 34], each bone can freely perform SE(3) transformations, making it challenging to ensure the physical plausibility of the internal skeleton. In experiments, we observed that while the output mesh was correct, the skin weights and bone transformations were not physically plausible for each variable. Additionally, the obtained bone transformations could not be directly used for motion transfer because the shapes, sizes, and part proportions between the reference object and the target object could differ significantly. Directly applying the same SE(3) transformation to corresponding parts would be unreasonable and would not ensure pose similarity. Therefore, in the second stage of reconstruction, we use kinematic chain skeletons and apply heat diffusion to obtain skinning weights, which narrow the deformation space of the skeleton, thereby ensuring the plausibility of the internal structure. This method uses forward kinematics to transfer pre-frame bone angles and bone scales, ensuring accurate motion (pose) transfer.
A.2 Video Results
We include an extensive set of 4D generation results in video format in Google Drive: link. There, we demonstrate 4D generation results across different species with diverse motions and compare the performance of pose transfer with 3D-CoreNet Song et al. (2021), X-DualNet Song et al. (2023), 4Dfy, SVD+Stag4D, DMT+Stag4D and Animate124 Zhao et al. (2023). Since the video files exceed 1000 MB, we are unable to upload them directly as a zip file. Therefore, we have used an anonymous Google Drive link.
Table 5: Prompts for User Study of 4D Generation.
# Prompts
1 A pig running like a rabbit
2 A pig attacking with its head
3 King Kong dancing hip hop
4 A pig dies and fall down
5 A pig jumping high and landing
6 A pig walking back
7 A pig jumptorting like a deer
8 Kung Fu Panda dancing hip hop
9 A penguin dancing hip hop
10 King Kong drunk walking
11 Kung Fu panda drunk walking
12 A penguin drunk walking
13 A pig attacking using its hind legs
14 A pig torting like a deer
15 A pig run like a tiger
16 King Kong hands up
17 Kung Fu panda hands up
18 A pig head down and drinking water
19 A pig stands and attacks with its left hand
20 A pig jumps on a box like a dog
21 A jelly man dancing hip hop
22 A Venom dancing hip hop
23 A unicorn attacking using its hind legs
24 A unicorn torting like a deer
25 A elephant run like a rabbit
26 A jelly man drunk walking
27 A Venom drunk walking
28 A rhino runs like a rabbit
29 A elephant stands and attack
30 A triceratops attack with its right leg
A.3 Dataset
A.3.1 Animal
Davis-Camel provides a real animal video in BADJA Biggs et al. (2018) with 2D keypoints and mask annotations, derived from the DAVIS video segmentation dataset Perazzi et al. (2016) and online stock footage. We extract reference motion from the reconstructed mesh sequence and transfer it to other identities.
PlanetZoo includes RGB synthetic videos of different animals with around $100$ frames each. PlanetZoo covers a 180-degree visual field captured by a moving camera to allow better evaluation of 3D reconstruction when imaging parameters must also be dynamically estimated due to the moving camera, and over a large visual field. In addition, following BADJA Biggs et al. (2018), we also provide 2D key point annotations.
DeformingThings4D is a synthetic dataset containing 1,972 animation sequences containing 31 categories of both humanoids and animals. Each sequence consists of 40 to 120 frames of motion animation. In this dataset, the first frame is the canonical frame, and its triangle mesh is given. From the 2nd to the last frame, the 3D offsets of the mesh vertices are provided, and we export the triangle meshes for all these frames. We use the motions of these animal mesh sequences as pose references and transfer the pose to different identities.
A.3.2 Human
EverybodyDanceNow consists of full-body videos of five human subjects. We use these monocular videos to generate human motions and transfer them to other identities.
DeformingThings4D also contains humanoid examples of dynamic mesh, as mentioned before. We use the motions of these sequences as pose references and transfer the pose to different identities.
A.3.3 Self-collected in-the-wild data
These in-the-wild data are collected from online resources, like the video of HipHop dancing from Youtube.
A.4 Implementation Details
For Davis-Camel, PlanetZoo, EverybodyDanceNow, and those self-collected data without ground truth mesh, we first train the Dual-Phase 4D Reconstruction Module on 2 NVIDIA L40S GPUs with batch size 16 for 10 epochs with a learning rate of $0.0001$. MagicPose4D uses Zero123 Liu et al. (2023) and SF3D Boss et al. (2024) as the image-to-3D model for 4D reconstruction. We then transfer the per-frame motion from Phase 2 of 4D Reconstruction to the target object with the Cross-Category Motion Transfer Module. The motion transfer is not learning-based and does not require any trainable parameters, which makes MagicPose4D generalize well to unseen identities and reference motions. For all other data with ground truth mesh available, we directly train the second phase of the 4D Reconstruction Module and then transfer the motions.
A.5 User Study
We provide a user study for comparison between MagicPose4D and previous works Song et al. (2021; 2023) on motion transfer. We asked 50 lay participants from Prolific, an online platform for user studies, to rate the quality of eleven in-the-wild retargeted mesh sequences from 3D-CoreNet Song et al. (2021), X-DualNet Song et al. (2023) and MagicPose4D on a scale of 0(low) to 5(high). The participants are paid with an hourly rate of 16 USD. In each mesh sequence comparison, we collect target identity mesh, reference motion mesh sequence, and retargeting results. We visualize them side-by-side. The retargeting results from different methods are anonymized as A, B, C, and the order is randomized. We provide video visualization of these comparisons (For each video, from left to right: Reference Motion; MagicPose4D; 3D-CoreNet; X-DualNet) in the Google Drive. Criteria for judgment: 1) The generated motion should match the reference pose mesh sequence. 2) The identity of the transferred mesh should match the identity reference. 3) The generated mesh sequence should be consistent and smooth. From the results presented in Tab. 3, we conclude that MagicPose4D provides the most satisfying generation of mesh sequences.
Similarly, we provide a user study for comparison with 4Dfy, DMT+Stag4D, and SVD+Stag4D. In each comparison experiment, we collect reference images, text descriptions of motion, mesh reference motion sequences, and generated videos from both methods. For 4Dfy we follow the official code and instructions, for the other two, we first use SVD to generate a video via text prompt or use DMT to do video-level motion transfer and then use Stag4D to do reconstruction from the generated video. For MagicPose4D, we use the reference image to generate the target mesh with an Image-to-3D model and transfer the motion of the reference mesh sequence. We compared the rendered videos from Animate124 to videos of generated mesh from MagicPose4D in different viewpoints. Criteria for judgment: 1) The generation’s identity should match the reference image. 2) The generated mesh sequence should be consistent and smooth. From the results presented in Tab. 1, we conclude that MagicPose4D provides the most satisfying visualizations.
Furthermore, we conduct a user study for comparison with STAG4D Zeng et al. (2024), 4Dfy Bahmani et al. (2024b), on 4D generation of articulated objects. We asked 50 participants from Prolific to rate the quality of nine sequences (3 objects from 3 views) on a scale of 0(low) to 5(high). In each sequence comparison, we collect reference images, text descriptions of motion sequences, and video results. We visualize them side-by-side. The video results from different methods are anonymized as A, B, and C, and the order is randomized. The Criteria for Judgment: 1) The overall generation quality - completeness; smoothness; consistency; etc. 2) The motion of the object in the video should match the text description. 3) The appearance of the object in the video should match the reference image. The settings for the participants are the same. We provide video visualization of these comparisons in the Google Drive.
@@ A.5 User Study
Table 6: User study of MagicPose4D on 4D Generation compared to Animate124 Zhao et al. (2023). The settings are the same as the study on Motion Transfer.
@@
A.6 2D Losses and Regularization
The 3D loss is described in the main article. Here we introduce the 2D losses and regularization terms. The 2D losses are similar to those in existing differentiable rendering pipelines Yang et al. (2021b; 2022b); Zhang et al. (2024c). We define $\mathbf{S^{t}}$,$\mathbf{I^{t}}$, $\mathbf{F}^{2\text{D},t}$ as the silhouette, input image, and optical flow of the input image, and their corresponding rendered counterparts as $\mathbf{\tilde{S}^{t}}$,$\mathbf{\tilde{I}^{t}}$, $\mathbf{\tilde{F}}^{2\text{D},t}$. The following losses ensure the fitting between rendered and original:
$\mathcal{L}_{\text{silhouette}}=\sum_{\mathbf{x}^{t}}\left\|\mathbf{s(x}^{t}\mathbf{)}-\mathbf{\hat{s}(x}^{t}\mathbf{)}\right\|_{2},$ (4)
$\mathcal{L}_{\text{optical flow}}=\sigma\left\|(\mathbf{\tilde{F}^{2\text{D},t}})-(\mathbf{F}^{2\text{D,t}})\right\|_{2}^{2},$ (5)
$\mathcal{L}_{\text{texture}}=\left\|\mathbf{\tilde{I}^{t}}-\mathbf{I^{t}}\right\|_{1},$ (6)
$\mathcal{L}_{\text{perceptual}}=\text{pdist}(\mathbf{\tilde{I}^{t}},\mathbf{I}^{t}),$ (7)
where pdist(,) is the perceptual distance Zhang et al. (2018). Also, we leverage three regularization terms: (1) Dynamic Rigidity term, which is introduced in paper.Zhang et al. (2024c). (2) Symmetry term, which encourages the canonical mesh to be symmetric.
$\mathcal{L}_{\text{symm}}=\sum_{i}\min_{j}\left\|v_{j}-\phi\left(v_{i}{}^{symm}\right)\right\|^{2},$ (8)
where, $v_{j}$ is the vertex j in the one side of the symmetry plane, and $v_{i}{}^{symm}$ is the vertex from the other side. $\phi$ is the reflection operation w.r.t to the symmetry plane. (3) Laplacian smoothing: We apply laplacian smoothing to generate smooth mesh surfaces.
$\mathcal{L}_{\text{shape}}=\left\|\mathbf{X}_{i}^{0}-\frac{1}{\left|N_{i}\right|}\sum_{j\in N_{i}}\mathbf{X}_{j}^{0}\right\|^{2},$ (9)
where, $\mathbf{X}_{i}^{0}$ is coordinates of vertex $i$ in canonical space. And $N$ is the number of vertices
Appendix B Broader Social Impacts
The proposed MagicPose4D for motion transfer offers extensive applications, enhancing communication in digital environments by enabling more effective self-expression through avatars or digital characters. Additionally, MagicPose4D has the potential to revolutionize the entertainment and media production industries by facilitating the creation of more lifelike and expressive characters in movies, video games, and animations.
However, this technology also presents potential negative social impacts. Privacy concerns arise from the unauthorized creation of realistic animations of individuals, and the technology could facilitate the spread of misinformation through deepfakes. Risks include job displacement in fields like acting and modeling, psychological effects from the blurring of reality and virtual experiences, and the exploitation of the technology for unethical purposes. Furthermore, cultural insensitivity, security threats, and the misuse of realistic animal animations could have broader societal implications. Addressing these issues requires robust ethical guidelines, legal frameworks, and technological safeguards to ensure responsible use and mitigate harm.